consider the simplest macro model with demand-determined output. the equations of the model are : C= 150+0.8YD, I = 400, G = 700, T = 0.2Y, X = 130, and IM = 0.14Y. Total autonomous expenditure in this month are?
1
Expert's answer
2019-07-05T09:11:02-0400

aggregate expenditure =C0+С *Y +I0+G0+X0-IM

autonomous expenditure=C0+I0+G0+X0

autonomous expenditure = 150+400+700+130=1380


answer: 1380